OpenStudy (anonymous):

A crafts store bought 27 1/4 ounces of potpourri. The shop divides the potpourri into 2 1/2 ounce bags and sells each bag for $4.25. a. How many full 2 1/2 ounce bags of potpourri can the shop make? Show your work. b. If all the bags are sold, how much money will the shop earn? and can u show your work because i want to know howto get the answer :)

OpenStudy (anonymous):

@AskMeLotss @Vocaloid

OpenStudy (welshfella):

Number of bags that can be made = 27 1/4 / 2 1/2

OpenStudy (anonymous):

ho choco sorry, I was doing something

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

(a) 27 1/4 (total bought) 2 1/2 (total of each bag used) To find how many 2 1/2 oz. bags can be made you need to divide 27 1/4 by 2 1/2 (27 1/4) / (2 1/2) (I am going to change the mixed number into improper fractions) (109/4) / (5/2) 109/4 x 2/5 = 109/10 = 10 9/10 ANSWER : 10 full size bags can be made. (b)10 bags sold at $ 4.25 each.....10 x 4.25 = $42.50 ANSWER : If all 10 bags are sold, the craft store will make $ 42.50.

OpenStudy (welshfella):

do do this convert the mixed fractions to improper fractions 27 1/4 = 4* 27 + 1 -------- = 109/4 4 2 1/2 = 5 /2 so we have 109/4 divided by 5/2 to do this invert the second fraction and multiply so it's 109/4 * 2 / 5

OpenStudy (anonymous):

ok thank you so much :)

OpenStudy (anonymous):

your welcome :D